What is the kinetic energy in kilojoules of a space station weighing 8042 kg and moving at a velocity of 2.603e+4 km/hr?
1
Expert's answer
2020-12-14T14:46:55-0500

Kinetic energy=?

Mass= 8042Kg

Velocity= 2.603e+4Km/h = 2603m/s (by conversion)

Kinetic energy = 1/2mv2

Kinetic energy = 1/2 x 8042x 20632

Kinetic energy = 1.71x1010J

Kinetic energy = 1.71x107Kj